### v2

# Standard Language Ideology in AI-Generated Language

Abstract
Standard language ideology is reflected and reinforced in language generated by large language models (LLMs). We present a faceted taxonomy of open problems that illustrate how standard language ideology manifests in AI-generated language, alongside implications for minoritized language communities and society more broadly. We introduce the concept of standard AI-generated language ideology, a process through which LLMs position ”standard” languages—and particularly Standard American English (SAE)—as the linguistic default, reinforcing the perception that SAE is the most ”appropriate” language. We then discuss ongoing tensions around what constitutes desirable system behavior, as well as advantages and drawbacks of generative AI tools attempting, or refusing, to imitate different English language varieties. Rather than prescribing narrow technical fixes, we offer three recommendations for researchers, practitioners, and funders that focus on shifting structural conditions and supporting more emancipatory outcomes for diverse language communities.
1. Introduction
Tools built on large language models (LLMs), such as ChatGPT, promise immense benefits to their users; however, those benefits are not distributed equally across speakers of different language varieties. Rather, ChatGPT—and other popular generative AI language technologies—reflect standard language ideology, which reinforces a hierarchy between language varieties and affirms ”standard” varieties (particularly ”standard” American English) as the most appropriate or legitimate form of communication. In this paper, we illustrate how and in what ways standard language ideology manifests in AI-generated language and present a novel taxonomy of open problems regarding standard language ideology in AI-generated language. Although concurrent work has surveyed challenges in natural language processing (NLP) for different languages and dialects (Joshi2025, ), this is (to our knowledge) the first to directly discuss the effects of language ideology in connection with LLMs.
We first ground our taxonomy in a literature review of research on standard language ideology, linguistic bias, and linguistic discrimination in LLMs (Section 2). This process was guided by our multidisciplinary team of researchers including linguists, computer scientists, and social scientists. We then discuss our taxonomy, which categorizes key open problems in this space and highlights their societal implications (Section 4). Finally, we discuss recommendations for developing more equitable language technologies (Section 5).
The taxonomy serves two primary purposes. First, it provides a conceptual framework for analyzing how linguistic biases in AI-generated language reflect and reinforce broader systems of power. This helps researchers identify the underlying language ideologies that shape AI development. Second, it functions as a diagnostic tool by linking those structural insights to concrete patterns in AI behavior. In this way, the framework supplies a theoretical lens for understanding why these biases arise, while the diagnostic function reveals how they materialize in practice.
Through grounding in sociolinguistic theory, we connect documented harm from AI-generated language to established research on language ideology. This connection allows us to show how well-documented hierarchies in language—including beliefs around which languages are considered ”standard” or ”better” and linguistic discrimination for speakers of minoritized varieties—are systematically reproduced in generative AI. By situating these dynamics within the framework of standard language ideology, we argue that such issues are not merely technical problems to be solved through engineering fixes, but structural problems that demand critical engagement with power and ideology. Accordingly, our recommendations aim not to mitigate particular harms, but to challenge and reshape the conditions that produce them.
2. Background
2.1. Standard language ideology
Standard language ideology is a construct that reinforces a hierarchy between language varieties. As defined by (lippigreen1994, ), it is an imposed “bias toward an abstracted, idealized, homogenous spoken language” whose goal is the “suppression of variation.” In other words, it is a common and false belief that some language varieties—usually those used by communities with more social prestige—are “better” or “more complex” than others. Linguistically, however, all language varieties are equally valid. The idea that certain languages are “better” than others ignores the fact that all language varieties are equally capable of expression (Hill2008, ). In fact, “standard” language is not spoken by any real community, but is an abstracted variety that can only be defined in contrast to the speech of marginalized communities.
Despite this, certain language varieties have been institutionally privileged as more “standard” and viewed as more “appropriate” than others. This privileged status is linked to the association with people in power that the language variety holds. As economic globalization has cast English as a lingua franca, or the common language adopted by speakers of different languages (Michaud2015EnglishAA, ), discrimination against different dialects of English has become intertwined with questions of discrimination against people of different nationalities and ethnicities (Crystal2003, ; Phillipson2010-nx, ; tietze2004spreading, ). The privileged position of English is a result of the social and historical power that English speakers have had throughout history, especially as it relates to colonization across many parts of the world. The spread of English as the default, and the increasing imposition of its standard varieties worldwide, has been framed as “linguistic imperialism” that threatens other languages and language varieties (Jenkins2013-kc, ; pennycook2004, ).
Research on linguistic discrimination details numerous cases of discrimination against speakers of non-standard varieties. In the United States, “Standard” American English (SAE) is considered the dominant language variety and reflects an abstracted collection of linguistic norms of middle-class, white men who have held disproportionate levels of power within the country (Baker-Bell2020-hf, ; Hill2008, ; Rosa_Flores_2017, ). Other language varieties have been devalued through their institutional subordination (Baker-Bell2020-hf, ; Hill2008, ; lippigreen1994, ). African American English (AAE) is one such language variety, among many others (e.g. Irish English, (moore2024, ); Indian English, (kutlu2020negative, ); Chicano English, (guerrero2014you, )). For example, using AAE has been linked to being denied housing due to “sounding Black” (Makoni2002-wb, ; Wright2023, ). Even when using ”standard” varieties such as SAE, marginalized people can be subjected to linguistic bias (samy2007critical, ). Though linguistic discrimination in English is most well-studied, similar discrimination occurs with varieties of other languages, such as Spanish (pascual, ; Romn2023, ) and German (Adler2019, ). Because language itself is linked to identity and power (crenshaw1995critical, ), language ideologies shape both social and linguistic structures (Woolard2020, ), in ways that can reinforce social identities and existing power imbalances.
By misleading people to believe that some languages and language varieties are better than others, standard language ideology can perpetuate harmful patterns of linguistic discrimination and the oppression of speakers of “non-standard” varieties. Linguistic discrimination can often serve as a proxy for other forms of discrimination along lines of race, gender, nationality, or class, among others (Drodowicz2024, ). This discrimination resulting from the promotion of language hierarchy can be subtle under perceptions of benevolence, such as encouraging others to speak more “appropriately,” or more obvious bigotry, such as associating certain ways of speaking with lacking intelligence (baugh2005linguistic, ). Relatedly, standard language expectations dictate access to social capital through means such as education, employment, or public office (Wiley1996, ). Given this, those who speak more closely to the standard varieties benefit from better access to such resources and opportunities.
2.2. Standard language ideology in large language models
Standard language ideology is reflected in LLMs in two key ways: performance discrepancies between language varieties and biases or stereotypes in outputs. Research on language models preceding the release of ChatGPT has highlighted ways in which these models perform worse for speakers of certain English language varieties. Several studies examine performance discrepancies related to AAE. For exmaple, language models perform worse for AAE on tasks including text generation, sentiment analysis, and parsing (Blodgett2018, ; field-etal-2021-survey, ; Groenwold2020, ). (thylstrupDetectingDirtToxicity2020, ) discuss how content moderation tools fail to capture the semantic richness of AAE, such as by making blanket assumptions about the complex semantics of reclaimed slurs. (Lin2024OneLM, ; ziems-etal-2022-value, ) measure performance on different NLP tasks for queries rewritten in AAE and find that tested models’ performance worsens on the rewritten queries. Furthermore, (deas-etal-2023-evaluation, ) discuss how language models including GPT-3, ChatGPT, and GPT-4 have higher perplexity for AAE, interpreted as greater difficulty in understanding AAE, and have trouble producing natural-sounding and semantically accurate AAE text. Outside of AAE, (yong-etal-2023-prompting, ) find inconsistent performance for ChatGPT’s production of Singlish [^1] Singlish is an English-based creole spoken in Singapore. and code-switched text involving Southeast Asian languages, finding that even when generated text was perceived as natural (as though produced by a native speaker), it sometimes contained “semantic inaccuracies…discernible by native speakers.”
In addition to issues of poorer performance for minoritized language communities, language models can also advance stereotypes regarding speakers of a particular language variety (hofmann2024dialect, ; nee2022, ). (finch2025findingvoiceevaluatingafrican, ) note issues in language models’ production of AAE, finding that AAE speakers prefer chatbots speaking ”mainstream” or ”standard” English varieties over AAE-speaking ones. (hofmann2024dialect, ) illustrate that language models produce text with harmful stereotypes about speakers of AAE through “dialect prejudice.’
At a higher level and illustrating the hierarchy of English language varieties in LLMs, in a study comparing eight non-standard varieties of English to Standard American and British English, (fleisig2024, ) found that language model outputs tend to default to standard varieties, particularly SAE, even in response to non-standard varieties. Responses to the eight non-standard varieties all had over 84% reduction in linguistic features from that variety, with over 96% reduction for Jamaican, Singaporean, Scottish, and African-American English. Meanwhile, outputs default to SAE orthography. Surveying native speakers of each variety to review AI-generated outputs, they also found that outputs to non-”standard” varieties consistently exhibit issues including lack of comprehension (10% worse compared to ”standard” varieties) and increased stereotyping (16% worse), condescension (12% worse), and demeaning content (22% worse). These studies illustrate that ChatGPT perpetuates linguistic discrimination against speakers of non-“standard” varieties in ways that further standard language ideology.
Performance discrepancies and stereotypes present in LLMs are linked to the language data that underpins these technologies. English is the default language for data powering large LLMs, a consequence of LLMs relying on language data from the Internet, particularly from the Common Crawl. Internet data overwhelmingly represents English, and Internet use varies due to social factors. An estimated 60% of all language content on the Internet is in English, despite only  17% of people speaking English globally (nee2022, ). Meanwhile, 88% of languages have “exceptionally limited resources” in digital spaces (joshi-etal-2020-state, ). Non-standard varieties of English are disproportionately removed from scraped Internet data when scraped pages are filtered using blocklists (dodge-etal-2021-documenting, ). Even if languages are well-represented in digital corpora, certain perspectives are over- or under- represented. For example, on Reddit, users are 67% male and 70% White, resulting in potential reinforcement of White, male perspectives ((bender2019benderrule, ), citing (barthel2016reddit, )). Other perspectives may be actively targeted, harassed, cyberbullied, and otherwise marginalized online. Research exposes growing concerns of online gender-based harassment with clear links between online spaces and misogyny, which can be further amplified for Black and brown women (ging2018, ; duggan2017online, ; rights2014violence, ). This harassment and the resulting safety and mental health implications can result in decreased use of online spaces. Meanwhile, although Black people have historically been overrepresented on Twitter compared to other demographics in the general US population (brock2012, ), Black Tweets are still often considered “inappropriate” and are more often inaccurately flagged as hateful by automatic hate speech detection tools (davidson-etal-2019-racial, ). Beyond being inaccurate, this disproportionately censors Black speakers.
Research on language models treats English—particularly SAE—as the status quo, with work on other languages “often considered ‘language specific’ and thus reviewed as less important than equivalent work on English” (bender2019benderrule, ). This default of SAE partly reflects the people who hold disproportionate power in these spaces. Leading industry developers of LLMs, such as Open AI, Google, Microsoft, and Meta, are largely headquartered in the United States, with employees and corporate leadership that skew White and male (omara2022why, ). Demographics of AI researchers in industry and academia are likewise skewed (west2019discriminating, ), with NLP researchers from Latin America, Africa, the Middle East, and Southeast Asia particularly underrepresented (rungta-etal-2022-geographic, ). While the demographics of those developing and leading NLP research and companies do not directly result in the reinforcement of SAE at the level of language models, this lack of diversity in the technologies’ developers and managers can lead to similar thinking in ways that miss potential biases the technology can have (cowgill2020biased, ) with a default towards an (inequitable) status quo.
Taken together, research illustrates the ways that standard language ideology extends to LLMs and AI-generated language, including through poorer performance and greater stereotyping for minoritized language varieties. Meanwhile, ”standard” English varieties—and particularly SAE—remains the default output.
3. Methodology
The taxonomy introduces a conceptual framework by synthesizing empirical evidence with insights from linguistic theory, sociolinguistics, and applied linguistics. It identifies and organizes a set of open problems that illustrate how standard language ideology is reproduced in generative AI outputs, alongside their broader societal implications. To identify open problems we conducted a review of large-scale studies on linguistic bias in LLMs, alongside a literature review on standard language ideology and linguistic bias. This linguistic background is essential for understanding the broader and longer-term societal implications of language technologies. While LLMs are a recent technological development, the ideological forces they reproduce have been extensively studied in prior research on language and power. Our taxonomy builds on this to surface how those same dynamics manifest in AI. After identifying relevant problems, we categorized and refined them through collaborative, interdisciplinary analysis among our multidisciplinary team. Each identified problem was linked to broader societal implications, with particular attention to harms experienced by minoritized language communities. We then iteratively refined the categories to ensure conceptual clarity and comprehensiveness. By grounding the taxonomy in both empirical studies of LLM behavior and critical sociolinguistic and AI ethics literature, this paper offers researchers and practitioners with a structured framework for understanding and addressing the reproduction of linguistic hierarchies in generative AI systems.
4. Taxonomy of open problems regarding standard language ideology in AI-generated language
We introduce a faceted taxonomy of open problems arising from reproduction of standard language ideology in AI-generated language technologies, alongside implications for different language communities globally (see Table 1). This taxonomy is structured around recurring trends identified in existing studies of linguistic bias in generative AI, with each facet linking a distinct problem to its broader societal consequences. In this section, we explore each open problem and its societal implications in detail.
Open problem Societal implications
The default production of “standard” language varieties reinforces “correct” ways of communicating. Subtly reaffirms language hierarchies and can reinforce or amplify people’s internal linguistic biases
Outputs have lower quality of service for minoritized varieties. Speakers of minoritized varieties experience more difficulty in using language models than speakers of “standard” varieties Speakers of minoritized varieties practice “digital code switching” to achieve better quality of service, with potential psychological tolls
Producing minoritized varieties can result in stereotyping of languages. Inaccurate reflection of the linguistic features of the particular variety Reinforcement of limiting or harmful stereotypes about minoritized language communities
Producing minoritized varieties can result in appropriation and/or manipulation. Appropriating the speech of minoritized groups decontextualize community norms and erases the history of the group’s struggles Non-native speakers of minoritized varieties prompting and using minoritized language varieties can result in “linguistic minstrelsy” or “figurative blackface” Facilitating production of minoritized varieties lowers the barrier for bad actors’ infiltration or manipulation of minoritized communities
Preventing outputs in minoritized varieties can result in limited quality of service and erasure. Lower quality of service from being unable to fully interact with the model in one’s native variety Erasure of minoritized language varieties and reinforcement of language hierarchies
Table 1. Taxonomy of harms regarding standard language ideology in AI-generated language.
The default production of “standard” language varieties reinforces “correct” ways of communicating. At a high level, generative AI technologies default to “standard” language varieties in their responses, particularly SAE, as illustrated by (fleisig2024, ). This subtly reinforces the idea that the “correct” or “more appropriate” way of communicating uses “standard” languages, and particularly SAE. This can have a cascading effect that impacts people’s own linguistic biases and perceptions of language hierarchies. Lessons can be learned from the dominant position that SAE has held in educational settings in the United States. Despite the prioritization of SAE in classrooms often being seen by teachers as a means of unifying students through a common language (delpit1992, ; Smitherman1986-fd, ), it has negative implications for minoritized students. For example, Black students who speak AAE feel compelled to adopt SAE and dominant ideologies, which then impact their identity expression, academic achievement, and self-perception (lyn2022, ).
Additionally, due to automation bias—a psychological tendency to over-rely on automation that can result in complacency regarding automation outputs (goddard2012automation, )—humans can have greater trust and overconfidence in AI (horowitz2024, ). If LLM outputs that prioritize SAE or disparage other dialects are seen as particularly trustworthy, they could exacerbate the idea that “standard” languages are more “correct” ways of speaking.
Outputs have lower quality of service for minoritized varieties. Generative AI technologies can produce lower-quality responses to speakers of minoritized language varieties. At a high level, these tools may not comprehend the user’s prompt as well, and result in an incorrect or unhelpful response (see Appendix A for an example). Relatedly, generative AI tools can more often inaccurately determine that inputs of certain minoritized language varieties are hateful or offensive speech, resulting in refusal to respond (davidson-etal-2019-racial, ).
These behaviors result in speakers of minoritized varieties experiencing more difficulty in using language models than speakers of “standard” varieties, which is a form of allocational harm (crawford2017, ). Furthermore, if users of generative AI tools are aware that using a minoritized variety results in poorer performance, they may be incentivized to use a “standard” variety instead, which can reinforce the stigmatization of non-standard varieties and result in “digital code switching.” Code switching, which results from a pressure minoritized speakers can feel to conform to more “standard” language, causes psychological tolls (mccluney2019, ). While additional research is needed, these psychological tolls could extend to the digital sphere.
Producing minoritized varieties can result in stereotyping of language use. Generative AI technologies may be prompted by the user to produce minoritized language varieties, or may be designed to respond to the particular language variety of the input. This may not always be a source of harm, particularly if users of that particular variety are asking for responses back in their language variety. However, responses in minoritized varieties can disproportionately carry or convey stereotypes or demeaning content.
In response to inputs of minoritized languages, models prompted to imitate the input variety may produce a stereotyped version of that language variety that does not accurately reflect the range of linguistic features used by that community. This behavior perpetuates the association between speaking a non-standard variety and stereotypical traits of that speaker community, which often serves as a covert form of racism, xenophobia, or other widespread harms ((rickford2016language, ); see also Section 2.1). The outputs could also convey demeaning content related to speakers of that language community. Examining ChatGPT outputs that imitated the language variety of the inputs, (fleisig2024, ) found that native speakers of minoritized varieties rated outputs as more often conveying demeaning content than responses to standard English varieties.
Producing minoritized varieties can result in appropriation and/or manipulation. Even if a language model correctly produces text that reflects the grammar of that minoritized variety, there may be concerns over appropriation (particularly if non-native speakers of that variety are asking for and using this text). Because speaking a non-standard variety can carry covert prestige (i.e., social value associated with use of canonically minoritized language varieties (labov2006social, )), use of non-standard varieties by non-native speakers, particularly White speakers appropriating AAE, has been discussed as “linguistic minstrelsy” or “figurative blackface” (Bucholtz2011, ; eberhardt2015, ). Parallel concerns have been raised regarding the use of speech, music, and image generation tools to imitate people of color, sometimes called “digital blackface” (MatamorosFernndez2022, ; sobande2021, ; Holliday2023, ; feffer2023deepdrake, ). As LLMs improve at mimicking minoritized varieties, these models may increasingly contribute to similar harms.
In addition, use of uncredited language from marginalized groups by language models can further cultural appropriation and erase the linguistic history of speaker communities. A popular “Gen Z Translator” extension of ChatGPT [^2] https://chat.openai.com/g/g-AbhjZGbhY-gen-z-translator that claims to “transform any short text to be Gen Z slang filled” frequently produces AAE. However, rebranding expressions that emerged in Black communities as generational vernacular, as opposed to vernacular linked to ethnic and racial communities, erases the history and contexts of these expressions. It risks “appropriating Black culture and perpetuating racism as [speakers] take on Black speech without assuming Black Americans’ struggle” (chery2022black, ).
Producing minoritized varieties could also result in manipulation of those language communities by people outside of them. Readily available ways of producing text that imitates a non-standard variety could help agents to feign in-group membership for malicious purposes, such as the case of Russian misinformation bots posing as Black people online (jerkins2022, ).
Preventing outputs in minoritized varieties can result in limited quality of service and erasure. Preventing language models from producing anything but prestige or “standard” varieties of a language avoids the potential harms discussed regarding appropriation and manipulation. However, if speakers of a marginalized variety want the model to reply in that variety, then it may constitute a quality of service harm because, unlike speakers of the “standard” variety, they are unable to fully interact with the model in their native variety. This can also contribute to erasure of minoritized languages. Thus, preventing production of AI-generated text in minoritized languages could further reinforce language hierarchies and the false notion of “correct” ways of speaking.
4.1. Standard AI-generated language ideology
Our taxonomy outlines issues of standard language ideology in AI-generated language, illustrating how popular language models and associated technologies grant more power to “standard” language varieties while opening opportunities for harm to speakers of minoritized language varieties. At a high level, default production of certain varieties reinforces “correct” ways of communicating that can have ensuing psychological implications for speakers of minoritized varieties. It reinforces an idea of standard AI-generated language ideology that holds SAE as a linguistic default and reinforces a linguistic bias that using SAE is the most “appropriate” way of speaking. AI-generated language thus subtly reinforces the belief that some language varieties (i.e. SAE) are better than others.
This is further reinforced by AI-generated language outputs being of lower quality for inputs in marginalized varieties compared to inputs in “standard” varieties. Given this, native speakers of minoritized languages may be incentivized to provide inputs in “standard” varieties even if they are less fluent in them, a form of digital code switching. If users prompt outputs to be in the particular minoritized language variety of the input, outputs can disproportionately carry stereotypes, produce demeaning content, facilitate appropriation, or support manipulation. Taken together, these harms can impact access to resources or opportunities when generative AI tools are used as gateways or checkpoints (e.g., chatbots for scheduling healthcare appointments or educational tutors) and as they become increasingly integrated into daily life. It illustrates the powerful, yet subtle, ways in which Western hegemony, and particularly American hegemony, manifests in and through emerging AI technologies.
4.2. No clear way for popular language models to “win”
These issues prompt a question regarding what constitutes desirable model behavior when different potential behaviors may result in different harms. Our taxonomy illustrates the challenges that technologists face in attempts to identify appropriate behavior regarding LLMs and different language varieties. What is the “appropriate” way that language models—and ensuing tools producing AI-generated text or voices—should behave given the current technological context?
Central to these dilemmas is the fact that people use language to reflect aspects of their identity. For example, speakers may switch to a non-standard variety when speaking to someone to signal their shared membership in an in-group (auer2013, ; gumperz1982discourse, ). However, the “identity” of a language model is unclear: is it assumed to reflect that of its creators (often a White, Western, and male-dominated group)? Is it, instead, a neutral tool with a “view from nowhere,” if such a view is even possible? Does the range of appropriate language use depend on who created the model, whose data was used to train it, who profits from it, or other factors? Furthermore, these decisions are typically made by model creators and tech industry members, but they fundamentally depend on the nuanced norms of speaker communities that rarely have a say in how these models are designed.
For popular generative AI chatbots and voice tools developed and owned by large corporations (e.g., ChatGPT, Gemini, Voice Engine) there is no clear “correct” behavior. These technologies draw on data from the Internet (which is largely comprised of “standard” language varieties, including SAE), resulting in inherent issues and inequities, including reinforcement of standard language ideology. If a model is designed to automatically mimic the input variety unprompted, as opposed to defaulting to the language variety it is largely trained on (i.e. SAE), this could lead to issues outlined in the taxonomy, such as appropriation and manipulation. Instead, some may argue that the default production of a “standard” language is acceptable or even desirable. However, this line of thought has been documented to cause harm to minoritized language speakers in other contexts. For example, in education systems, prioritization of SAE results in psychological and academic impacts for Black students (lyn2022, ), while digital code switching may carry psychological tolls. The prioritization of SAE in ChatGPT and other popular generative AI technologies could result in harmful impacts at an even greater, global scale. Therefore, it is not clear what immediate approach should be taken for behavior of generative AI language technologies. This lack of clarity is concerning, particularly as those who hold decision-making power regarding the largest generative AI language technologies are their for-profit corporate owners with incentive structures tied to delivering value to shareholders.
5. Recommendations: Moving towards emancipatory outcomes
Instead of asking what the “appropriate” way is for generative AI models to behave, we proposal a more foundational question: how might the recognition of standard language ideology guide the development of generative AI models towards more emancipatory and socially just outcomes? This reframing is crucial because there is no universally ”correct” way for models to respond across all contexts (e.g., there is not a clear answer around what the default production of models should be), and a narrow focus on performance can obscure deeper social and political stakes that standard language ideology calls us to examine. Therefore, rather than aligning our recommendations to each individual open problem, we instead take a higher-level perspective to focus our recommendations on addressing the underlying power dynamics that shape AI systems, including questions of ownership, control, and community agency. We present three concrete actions for AI researchers, funders, and practitioners.
1. Expand evaluation metrics to include language variety harms. First, and most simply, the fact that language models perpetuate discrimination on the basis of language variety means that evaluation of language models for potential harms (e.g. “toxicity” evaluations) should expand to capture discrimination based on the user’s language variety. Measurement of harms due to standard language ideology is particularly important because language use often serves as a proxy for other harms–a way that racism, sexism, or xenophobia may fly under the radar.
We therefore recommend that AI developers and researchers do the following: (a) Incorporate language variety as a specific category in model evaluation frameworks alongside race, gender, and other demographic attributes; (b) develop benchmarks and metrics that assess whether models perpetuate harms (e.g., erasure of linguistic features, stereotyping, and condescension) against different minoritized language varieties; and (c) publicly share evaluation results to ensure transparency and accountability, while raising awareness about the risks of reinforcing linguistic discrimination and hierarchies through AI generated language. By incorporating language variety as a category for evaluating performance and ”toxicity” and being transparent about results, we can better capture and understand the implications of AI technologies on marginalized language communities globally.
2. Support speakers of different language communities in creating their own datasets and models. Issues surrounding appropriate responses to minoritized varieties, and when/if responding in the variety is appropriate, stem in part from the fact that these models are developed by organizations that are not composed of members of these speech communities. Instead, the development of language models by speaker communities, for speaker communities, helps to avoid this top-down imposition of linguistic norms (see, for example, (mahelona2023openai, ; nekoto-etal-2020-participatory, ), and the broader discussion of indigenous data sovereignty by (kukutai2016indigenous, )).
There are various examples of communities whereby speakers of minoritized language varieties are collecting data and developing language models. One example is Masakhane (orife2020masakhane, ), a grassroots community focused on NLP research “in African languages, for Africans, by Africans.” With over one thousand participants in their community across 30 African countries, they have developed translation results for approximately 40 African languages and supported open-source datasets and subsequent machine translation models to address under-representation of African languages in technology. Another example is AI4Bharat (kunchukuttan2020ai4bharat, ), a research lab at IIT Madras that ”is dedicated to advancing AI technology for Indian languages through open-source contributions.” The organization seeks to build language models and datasets for various Indian languages across India’s diverse linguistic landscape. We recommend additional support for these types of initiatives in which speakers of different communities are collecting data and developing models for their communities. In providing support to such initiatives, ensuring data collection is not extractive and providing funding for data collection efforts and people whose data is collected remains important.
There are some cases that go further by ensuring that people of their community have a say in how that data is used. Te Hiku Media – a Māori-owned and operated nonprofit based in New Zealand – focuses on preserving and promoting the Māori language (te reo Māori) and culture (james2020developing, ). In 2018, they launched a competition in which Māori speakers across New Zealand recorded over 300 hours of annotated audio. With this, they built AI tools for automatic speech recognition and pronunciation of te reo Māori. Te Hiku Media created various AI technologies and tools for Māori (e.g., automatic speech recognition, auto tagging, real time language feedback). They also established new data protection protocols and worked with Māori-led organizations to develop AI tools that support the community. Simultaneously, they worked with other Indigenous communities to leverage the Te Hiku language database to train language models for other Indigenous groups. The diversity of methods adopted by these organizations–cross-national collaboration, open-source contributions, and community-centered competitions–illustrates how unconventional methods of research can help to center the needs of non-“standard” speech communities. In 2024, a data sovereignty agreement between Microsoft and Te Tumu Paeroa (the office of the Māori trustee) was established, in which Te Tumu Paeroa, acting as the anchor tenant for Microsoft’s New Zealand-based data centers, ensured that Māori data would be stored locally and governed according to a Māori data sovereignty framework. This agreement reflects an institutional recognition and honoring of Indigenous data governance that supports community ownership.
We recommend that multilateral institutions, funders, and policymakers prioritize meaningful support—financial, technical, and infrastructural—for language communities’ own dataset and AI development efforts. Community-led development helps prevent the top-down imposition of linguistic norms and enables more contextually appropriate, culturally-informed systems. As the approaches we highlight demonstrate, these types of efforts are already underway and can be further enabled with targeted support. Crucially, supporting language communities also means respecting their preferences and boundaries. In some cases, communities may choose to practice refusal—opting not to share or use their data in the development of AI technologies by outside actors—or may prioritize data sovereignty above participation. Supporting diverse language communities means supporting these choices, including if and when they resist mainstream technological integration.
3. Innovate on and implement participatory processes that center marginalized communities, while rethinking language data practices to avoid extractive approaches. Given the dominance of a small set of LLMs, such as ChatGPT, there is also an onus for developers of these widely used models to improve model quality for speakers of different language varieties (not only rely on those communities to address linguistic data and performance gaps). In this case, we recommend that AI researchers and practitioners: (a) Engage minoritized language communities as equal partners throughout the AI development lifecycle, from data collection and model training to deployment and evaluation (b) support community ownership by developing language models by and for speaker communities that reflect the values, norms and linguistic practices of communities they serve; and (c) implement participatory governance structures that give communities the power to make decisions about how their language data is used and tools are deployed.
An example here is the partnership between OpenAI and the Icelandic government. While most of Iceland’s citizens speak English or another second language, its integration with the United States and Europe has put the country’s native tongue, Icelandic, at risk. With concerns that Icelandic cannot remain the default language in the face of digitalization, the Icelandic government partnered with OpenAI to use GPT-4 in the preservation effort of the Icelandic language (iceland, ). As part of this initiative, $\eth$eind ehf, a language technology company, assembled a team of 40 volunteers to train GPT-4 on proper Icelandic grammar and cultural knowledge. In this case, AI practitioners (with the help of the particular community) recognized that the model was not serving the language community and sought to better learn from the speakers to improve model performance for the language variety. However, OpenAI is benefiting from the engagement of the particular linguistic community and it is not clear to what extent (if at all) the linguistic community is sharing in the financial benefits over time. This initiative could go further by ensuring that Icelandic speakers are central actors in the development and management of generative AI tools that use and leverage their language and the work they provided to improve the models.
Despite the difficulty of implementing participatory practices in model design, setting “ambitious, yet reasonable goals” that help to move from one of the “extreme poles of transactional consulting and transformative empowerment” to the other (delgado2023, ) could help to improve power dynamics involving speakers of minoritized varieties. In addition, participatory AI practices include compensating them fairly for their contributions and respecting indigenous data sovereignty principles. While participatory processes may require additional time and resources, they are critical to create AI systems that empower rather than further marginalize communities. AI researchers and practitioners must rethink their approach to language data by recognizing what language data these models are (and are not) learning from, how they can better learn from minoritized language varieties in non-extractive ways, and how marginalized communities can be enabled as central actors in the development and ownership of generative AI tools.
We recommend that AI researchers and practitioners: (a) Audit existing language datasets to identify gaps in representation for different language varieties and be transparent about linguistic data representation in models and about what languages the models work better (and worse) for; (b) incorporate diverse language into training data, ideally in ways that build from recommendation 2; and (c) collaboratively develop mechanisms to compensate and credit communities whose language data is used in model training. As researchers strive to move from merely consulting language communities to empowering them regarding the use and management of their data (delgado2023, ), the examples we describe in each of these directions can serve as case studies for future collaboration.
These actions require a mindset shift towards prioritizing equity considerations and valuing linguistic diversity as a fundamental aspect to AI development, as well as securing additional resources. Although these actions may increase development costs and timelines, they are critical to addressing the harms perpetuated by current language models and creating more emancipatory technological futures.
6. Conclusion
In this paper we identified how standard language ideology is generated and perpetuated by LLMs that are powering some of the largest generative AI applications used today, resulting in “standard” AI-generated language in which SAE is the linguistic default and holds the linguistic bias that SAE is the most “appropriate” language. Our taxonomy of open problems regarding standard language ideology in AI-generated language captured the implications for different language communities. While tensions remain, including questions around what desirable system behavior looks like, we argue that the focus should instead be on moving towards emancipatory outcomes. We present three actions for AI researchers, practitioners, and funders: (1) expand evaluation metrics to include language variety harms; (2) support speakers of different language communities in creating their own datasets and models; and (3) innovate on and implement participatory processes that center marginalized communities, while rethinking language data practices to avoid extractive approaches.
7. Limitations and Ethical Considerations
This paper highlights the ways in which standard language ideology is present in AI-generated language, with implications for minoritized communities and society more broadly. While we believe that our taxonomy on open problems considers a range of issues, the list and associated implications are likely not exhaustive. A potential consequence of overlooking some work is narrowing the range of perspectives considered in future research based on the avenues we mention and the structure of our taxonomy. Nonetheless, it represents a first step in the development of a complete taxonomy and introduces critical questions to inform forward paths. Our discussion also centers largely on standard language ideology in English; the manifestation of standard language ideology in other languages is an important area for future work.
Appendix A Example response
Here, we provide an example of an incorrect and unhelpful response to a prompt in Indian English. Part of an input in Indian English stated: “I hope you have received my earlier letter. I would have liked to take up your kind invitation to act as resource person for the ELTC seminar, but for reasons already mentioned, I won’t be able to make it this time.”
When prompted to respond, ChatGPT’s response simply rephrased part of the input: “Thank you for your letter dated Feb. 16th. I appreciate your kind invitation to act as a resource person for the ELTC seminar. However, due to reasons mentioned in my previous letter, I regret to inform you that I won’t be able to attend this time.”